Find answers to the following. Write and indicate how you solved them.
 Is 4/ 5  equal to 16/20

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

4 / 5, 16 / 20
16 / 20 = (4 × 4) / (5 × 4)
= 4 / 5
∴4 / 5 = 16 / 20
Hence, 4 / 5 is equal to 16 / 20

​Thanks
Harshit Singh
3 years ago
Dear Student

4 / 5, 16 / 20
16 / 20 = (4 × 4) / (5 × 4)
= 4 / 5
∴4 / 5 = 16 / 20
Hence, 4 / 5 is equal to 16 / 20

​Thanks